OpenStudy (anonymous):

log{b} 1000=3 I know b=10 but I don't understand how they got there.

OpenStudy (anonymous):

\[\log_{b} 1000=3\]

OpenStudy (anonymous):

\[b^3=1000\] \[3\sqrt{b^3}= 3\sqrt{1000}\] b=10

Nnesha (nnesha):

well did you take 3rd root right ? you can write 1000 in terms of 3 exponents

Nnesha (nnesha):

x^3 =1000 ? what is x

OpenStudy (anonymous):

10

OpenStudy (anonymous):

What is the point of the 3's in front of the square root signs?

Nnesha (nnesha):

that's not 3 it's a 3rd rooot i guess like this \[\huge\rm \sqrt[3]{b^3}\]

OpenStudy (anonymous):

Right, I didn't know how to do that

Nnesha (nnesha):

you can convert 3rd to 1/3 exponent \[\huge\rm \sqrt[3]{b^3} = b^{3 \times \frac{ 1 }{ 3 }}\] so cancel out 3 exponent u have to take 3rd root :=)

Nnesha (nnesha):

\[\huge\rm \sqrt[3]{b^3} = b^{\cancel3 \times \frac{ 1 }{\cancel 3 }}=b\]

OpenStudy (anonymous):

Wouldn't that also mean that 1000 would be multiplied by 1/3?

Nnesha (nnesha):

yes so 1000 can be written as 10 to the 3rd power \[\huge\rm \sqrt[3]{10^3}\]

OpenStudy (anonymous):

See that's where I get confused, so the 1000 get simplified down to 10 because?

Nnesha (nnesha):

because it's easy to solve without calculator and it's legal :3

OpenStudy (anonymous):

-_- so because 10^3=1000 it gets simplified to 10^3 and since it's the 3rd root the exponent cancels out leaving b=10 ?

Nnesha (nnesha):

b to the 3rd power so you can either write 10 to the 3rd power or put 3rd root of 1000 into the calculator :=)

Nnesha (nnesha):

yep

OpenStudy (anonymous):

See when I enter that into my calc I get 94.86....

Nnesha (nnesha):

and it's helpful when you have to find an exponent given base for example 10^x = 1000

Nnesha (nnesha):

what did you enter ?

OpenStudy (freckles):

\[3 \sqrt{1000} \neq \sqrt[3]{1000}\]

OpenStudy (anonymous):

\[3\sqrt{1000}\]

OpenStudy (anonymous):

Yea see I don't know how to enter the latter

Nnesha (nnesha):

it's not 3 times square root of {1000} 3rd root like \[\sqrt[3]{1000}\]

Nnesha (nnesha):

ohh then convert 3rd root to 1/3 exponent \[\huge\rm \sqrt[3]{1000} = (1000)^\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

That works. =] Could you possibly help me with another problem?

Nnesha (nnesha):

gtg :3 sorry

OpenStudy (anonymous):

All's well Thank you for your help :)

Nnesha (nnesha):

my pleasure